OpenStudy (anonymous):

What is this quadratic equation written in standard form? x^2 - 10x - 6 = 2x + 1

Parth (parthkohli):

\( \color{Black}{\Rightarrow ax^2 + bx + c = 0 }\) This is the standard form.

OpenStudy (anonymous):

Group like terms, sucka.

OpenStudy (jit4won):

x^2 - 10x - 6 - 2x - 1= 0

OpenStudy (anonymous):

got it,thanks.

Parth (parthkohli):

Well if you bring like terms to one side THEN it'd be the standard form.

OpenStudy (anonymous):

x^2-12x-7=0

Parth (parthkohli):

Correct xD

OpenStudy (jit4won):

yeah

OpenStudy (anonymous):

GENIUS!

OpenStudy (jit4won):

who ? you ?